OpenStudy (anonymous):

Given the exponential equation 3^x = 27, what is the logarithmic form of the equation in base 10?

OpenStudy (anonymous):

I don't remember the formula to convert it or what to do with the base 10

OpenStudy (perl):

you don't need a logarithmic equation here $$ \Large 3^x = 27 \iff 3^x = 3^3 $$

OpenStudy (perl):

to use logs we have

OpenStudy (perl):

$$ \Large 3^x = 27 \iff x = \log_{~3} 27 $$

OpenStudy (perl):

you have to use change of base formula to make that base ten log

OpenStudy (anonymous):

log3(27 = log10(27) / log(10)3

OpenStudy (anonymous):

right? can't figure how to write in equation calculator thing haha

OpenStudy (loser66):

yup

OpenStudy (anonymous):

thanks both of you